OpenStudy (anonymous):

For which system of equations is (2, 2) a solution? A. –3x + 3y = 0 x + 6y = 10 B. –2x + 5y = –6 4x – 2y = 4 C. 5x – 2y = –6 3x – 4y = 2 D. 2x + 3y = 10 4x + 5y = 18

OpenStudy (johnweldon1993):

So you have an ordered pair (x , y) (2 , 2) so all you would do is plug in (2) for 'x' and (2) for 'y' in each equation you have...and see which one works...for example A) –3x + 3y = 0 x + 6y = 10 Plug in 2 and 2 for each 'x' and 'y' you see... –3(2) + 3(2) = 0 -6 + 6 = 0 so this part works... 0 = 0 and (2) + 6(2) = 10 2 + 12 = 10 14 = 10 This obviously isn't correct.....so A is not your answer...just do this for the other options and see which one works

OpenStudy (anonymous):

thanks

OpenStudy (johnweldon1993):

No problem!